OpenStudy (anonymous):

I GIVE MEDALS! What is the average velocity of atoms in 1.50 mol of neon (a monatomic gas) at 278 K? For m, use 0.01500 kg. Remember that R = 8.31 J/(molK). A. 9610 m/s B. 832 m/s C. 6.93 m/s D. 693,000 m/s

OpenStudy (lastdaywork):

See if this makes sense - http://en.wikipedia.org/wiki/Kinetic_theory#Temperature_and_kinetic_energy

OpenStudy (anonymous):

i tried looking at that before i asked. I dont think i understand like any of it, to be honest. please help cx @LastDayWork

OpenStudy (lastdaywork):

Use - PV = NRT PV = (2/3)K K = (1/2)mv^2 BTW what is "cx" ??

OpenStudy (anonymous):

You should use Maxwell-Boltzmann velocity distribution..

OpenStudy (anonymous):

\[v _{average} =\sqrt{\frac{ 8RT }{ \pi M }}\]